Please help!
1.) Find all zeros(both real and complex) of the function f(x)=x^4+4x^2-5


2.) Will sells used 4-wheelers and bicycles. He currently has 46 of these machines for sale. If there are 146 wheels on all of these machines, how many 4-wheelers and how many bicycles does Will have for sale?
Thank you!!!!

Answer by stanbon(57290)

Find all zeros(both real and complex) of the function f(x)=x^4+4x^2-5
=(x^2+5)(x^2-1)
=(x+isqrt5)(x-isqrt5)(x+1)(x-1)
Zeroes are 1,-1,isqrt5,-isqrt5
----------------------


2.) Will sells used 4-wheelers and bicycles. He currently has 46 of these machines for sale. If there are 146 wheels on all of these machines, how many 4-wheelers and how many bicycles does Will have for sale?
Let # of 4-wheelers be W and # of bicycles be B:
B+W=46
2B+4W=146
--------------
B=46-W
2(46-W)+ 4W=146
92-2W+4W=146
2W=54
W=27 (# of 4-Wheelers)
)B=46-27=19 (# of bicycles)
---------------
Cheers,
Stan H.